# Adaptive Energy Regularization for Autonomous Gait Transition and Energy-Efficient Quadruped Locomotion

### v1

@@ IV-A ANYmal-C Settings
We utilized the robot model and PPO training package in [4]. The system outputs the position command of the 12 joints in the next time step. The system inputs include the linear and angular velocities of the trunk, projected gravity in the robot frame, the commanded x-y velocities, the commanded yaw rate, each joint’s position and velocity, as well as the action at previous time step. The commanded y-velocity and yaw rate are fixed at zero here; only the commanded x-velocity will be set. The training episode will reset after 1000 time steps or if any part of the robot except its feet touches the floor.
The policy was trained on a flat ground with the coefficient of friction randomized between $[0.0,1.5]$. We also disturbed the mass of the robot with a uniform random value in $[-5.0,5.0]$ kg. A uniformly distributed noise was added to the observation. A random push with x-y velocity uniformly sampled between $[-1,1]$ m/s lasting 15 seconds was exerted on the robot. All these randomized domain parameters are renewed every time the training episode resets, except observation noise is resampled after every time step.
In motion rewards (3), during fixed velocity training (to get results in Figure 3) and when the $\hat{v}_{x}$ is large, the policy may fail to converge to a desirable tracking accuracy. Hence, we use $\sigma_{v}(\hat{v}_{x})=|\hat{v}_{x}|^{2}/3$ to overcome this difficulty and $\sigma_{\omega}$ is fixed at $0.25$. In energy rewards (4), the energy scaling constant $\sigma_{en}$ is fixed at $800$. To obtain the energy weight curve $\alpha(\hat{v}_{x})$ and the normalization index $Z(\hat{v}_{x})$, we trained fixed-velocity policies by setting $\hat{v}_{x}$ at $0.5$ to $2.0$ m/s, with $0.1$ common difference. For each fixed reference velocity, we trained $11$ policies using different $\alpha(\hat{v}_{x})$ values ranging from $0.5$ to $4.5$. Figure 3 summarizes the corresponding $\alpha(\hat{v}_{x})$ and $Z(\hat{v}_{x})$ for each reference velocity. These dense parameter pairs are exhibited to visualize the $\alpha$-$\hat{v}_{x}$ relation; in real experiments, only a few pairs of parameters at representative speeds are required.
Figure 3: $\alpha(\hat{v}_{x})$ and $Z(\hat{v}_{x})$ corresponding to each reference velocities. These two values are defined in (2). For each reference velocity, fixed-velocity trainings were conducted across various $\alpha(\hat{v}_{x})$, then we select the largest $\alpha(\hat{v}_{x})$ whose velocity tracking error is smaller than a threshold.
(a) Comparison of energy consumption
(b) Comparison of velocity tracking error
Figure 4: Ablation study in ANYmal-C simulation. Reference velocities are chosen from $0.5$ to $2.0$ m/s with $0.1$ common gap. Results of each velocity are obtained from $50$ independent runs. The solid line indicates the average and the shadow indicates the variance. Left: Energy consumption per unit moving distance under different reference velocities. We can see that adaptive energy regularization generates the most energy-efficient policy compared to fixed energy regularization. The $\alpha=0$ policy is neglected because it consumes an uncontrollable amount of energy. Right: Velocity tracking error under different reference velocities. Among all fixed energy regularization policies, only $\alpha=0.9$ has a comparable tracking error. However, the left figure shows that $\alpha=0.9$ policy is considerably less energy efficient.
IV-B Go1 Settings
@@
We mainly inherited the training methods released by [3]. Similar to [4], the system also outputs the position command of the 12 joints, but its input excludes the linear and angular velocities of the trunk. In addition, the inputs of the previous 30 time steps are also given to the RL system. Compared to ANYmal-C, we found that the energy reward $R_{en}$ alone is insufficient to regularize Go1’s behavior, which is likely due to the lighter weight compared to its motor power. Thus, following the settings in [3], we further add a fixed auxiliary reward $R_{aux}$ to (2) as an amendment to the normalization index $Z(\hat{v}_{x})$. This auxiliary reward is derived mainly from safety concerns, such as penalizing limb-ground collision, out-of-range joint position, and high frequency joint action. The details of $R_{aux}$ can be found on the project website. The training episode will reset after 1000 time steps or if the trunk touches the floor. Similar domain randomization in section IV-A is applied here. Compared to [3], we did not include any gait-related rewards.
We also discovered that curriculum technique is essential even for fixed velocity training. Given a reference velocity $\hat{v}_{x}$, the sampling range of x-velocity starts with $[-\min\{\hat{v}_{x},1\},\min\{\hat{v}_{x},1\}]$ m/s. the sampling range increases when the total reward achieves a certain threshold, and the maximal sampling range is set at $[-\hat{v}_{x}-0.1,\hat{v}_{x}+0.1]$ m/s. The fixed velocity training is considered a success if the x-velocity sampling range expanded to $[-\hat{v}_{x}-0.1,\hat{v}_{x}+0.1]$ m/s in the end and the tracking error for speed $\hat{v}_{x}$ is smaller than a threshold $\delta$.
In motion rewards (3), both $\sigma_{v}(\hat{v}_{x})$ and $\sigma_{\omega}$ were fixed at $0.25$. In energy rewards (4), the energy scaling constant $\sigma_{en}$ is fixed at $300$. We trained fixed-velocity policies by setting $\hat{v}_{x}$ from $0.5$ to $2.5$ m/s, with a 0.5 common difference. For each reference velocity, we trained eight policies using different $\alpha$ values ranging from $0.7$ to $2.1$, then fit the $\alpha(\hat{v}_{x})$ curve with the method in section III.
V Experiments
The experiments were designed to show the following after the legged robot was trained using the adaptive energy-regularized reward shown in Equation (2).
  • With adaptive $\alpha(\hat{v}_{x})$ and $Z(\hat{v}_{x})$, the legged robot automatically selects suitable behaviors to move with the reference velocity. We also demonstrate that the trained policy is deployable to a real quadruped robot.
  • If $\alpha(\hat{v}_{x})$ and $Z(\hat{v}_{x})$ are constants, the legged robot may not be able to find an energy-efficient walking policy for all target velocities.
  • Our method can generate a more energy-efficient walking policy compared to established baselines.
Figure 5: Three policies walking at $0.5$ m/s. Built-in MPC and walk-these-ways (WTW) [3] swing the legs to a redundant height, while our policy only makes necessary lifting of the leg. This demonstrates the energy-efficiency of our policy.
V-A One Policy with Different Gaits at Different Velocities
@@
To demonstrate natural emergence of efficient locomotion gaits, we evaluate the trained walking policy under different reference velocities. Fig. 2 demonstrates a trial run where the legged robot was commanded to move at $\hat{v}_{x}=0.5$, $1.0$ and then $2.0$ m/s. Each commanded velocity lasts for two seconds. We plot the gait recorded from the ANYmal-C simulation and show the snapshots for simulated ANYmal-C and real world Go1. We can see that our trained policy exhibits four-beat walking at low speed ($0.5$ m/s) by moving one leg each time in the order of right hind, right front, left hind and left front. At medium speed ($1.0$ m/s), the policy exhibits an intermediate gait between walking and trotting. At this gait, the right hind and left front legs move at the same time, while the right front and left hind legs have a displacement in their motions. At high speed ($1.0$ m/s), the trained policy exhibits a standard trotting gaits, where the right hind and left front legs move together, and the right front and left hind legs move together. This gait transition is endorsed by previous works [1, 21] that four-beat walking and trotting are respectively the most energy-efficient gaits under low and high speeds.
A closer look at the snapshots in Figure (2) uncovers that the swing ratio was also regularized. The legged robot only lifts up its leg to a height necessary to reach the reference velocity $\hat{v}_{x}$ to avoid wasting energy. At low speed, it only mildly lifts up its feet. As the reference velocity increases, the robot also lifts its feet higher, but only to a necessary height. Without considering energy-efficiency, the trained policy usually tends to lift the feet redundantly high. This will be argued in more detail in section V-C.
Finally, Figure 5 also showcases a successful hardware deployment of our policy. Videos can be found in our project website stated in the abstract.
V-B Ablation Studies with Fixed Energy Rewards
In ANYmal-C simulation, when the reference velocity is fixed at $1.0$ and $2.0$ m/s, we found that velocity tracking error was reasonably small when $\alpha$ was set as $3.0$ and $0.9$. Therefore, we run ablation experiments by fixing $\alpha$ at these values and $0.0$ to compare with varying $\alpha(\hat{v}_{x})$.
Figure 4(a) shows the energy consumption for adaptive $\alpha$, $\alpha=0.9$ and $\alpha=3.0$. The policy with $\alpha=0.0$ is dropped because it consumes multiple orders more energy. We observe that adaptive $\alpha$ reaches the lowest energy consumption. Figure 4(b) shows the velocity tracking error. Only the $\alpha=0.9$ policy has comparable tracking accuracy with adaptive $\alpha$, but Figure 4(a) shows that $\alpha=0.9$ policy consumes considerable more energy. These observations conclude the non-triviality of finding a constant energy reward scale $\alpha$ and elaborates the usefulness of $\alpha(\hat{v}_{x})$ varying with reference velocity.
V-C Comparison to Other Methods on Go1
We also compare our method with built-in MPC and walk-these-ways on real world Go1 robot. All policies are commanded to move at $0.5$ m/s. Such low velocity requires only mild movement of each leg for Go1. Figure 5 shows snapshots of each policy. It can be seen that both built-in MPC and walk-these-ways over-swing the legs, which squanders energy. Our policy swings the leg only to a necessary height, which can be visually recognized as the most energy-efficient.
VI Discussions
This paper focuses on developing energy-efficient locomotion strategies for quadruped robots, employing a simple yet effective reinforcement learning approach. However, there are some inherent limitations which provide avenues for future research.
VI-1 Limitations
The core limitation of the presented approach lies in the requirement for pre-running experiments to determine appropriate energy regularization weights. This necessity stems from our method’s reliance on empirical observations to calibrate the weights, which, while effective, does not afford the flexibility in a fully adaptive reinforcement learning system. Although the policy developed is applicable across different speeds post-training, it cannot be obtained with only one training for a new quadruped platform. Moreover, we only tested its generalizability across speeds. We did not verify the adaptive capability of energy rewards in different environments, which would be crucial for deploying these robots in real-world scenarios where they might encounter multiple operational challenges.
VI-2 Future extensions
Future research could address these limitations to develop methodologies for automatically tuning energy regularization weights within one single reinforcement learning training. This would enable the system to dynamically adjust its strategy in response to multi-task RL [33] or cross-embodiment settings [34, 35, 36].
Moreover, while this study concentrated on locomotion tasks, the underlying principle of leveraging energy efficiency to drive behavior selection holds broader potential. Future work could explore applying this energy-centric approach across different robotic tasks. For instance, manipulation and interaction tasks could also benefit from strategies prioritizing energy efficiency, potentially finding natural, efficient behaviors analogous to those observed in biological systems [37, 38]. Such a framework would align robotic systems more closely with sustainability principles and environmental consciousness.
VII Conclusions
@@
This paper presented a novel approach to energy-efficient locomotion in quadruped robots through the implementation of a simplified, energy-centric reward strategy within a reinforcement learning framework. Our method demonstrated that quadruped robots, specifically ANYmal-C and Unitree Go1, could autonomously develop and transition between various gaits across different velocities without relying on predefined gait patterns or intricate reward designs. The adaptive energy reward function, adjusted based on velocity, enabled these robots to select the most energy-efficient locomotion strategies naturally. Our policy showed energy-efficient behaviors and gait transitions in both simulation experiments (ANYmal-C) and hardware experiments (Go1). We also demonstrated the usefulness of adaptive energy regularization via ablation studies.

### v2

@@ IV-A Training Setup
We utilized the robot model and PPO training package in [4]. The system outputs the position command of the 12 joints in the next time step. The system inputs include the projected gravity in the robot frame, the commanded x velocities, the commanded yaw rate, each joint’s current position and velocity, as well as the action at previous time step. In addition, the inputs of the previous 30 time steps are also given to the RL system. The training episode will reset after 1000 time steps or if any part of the robot except its feet touches the floor.
IV-A1 Details of Reward Parameters
@@
In motion rewards (3), both $\sigma_{v}$ and $\sigma_{\omega}$ were fixed at $0.25$. In energy rewards (4), the energy scaling constants are fixed at $\sigma_{en,x}=1000$ and $\sigma_{en,z}=500$. As stated in section III, we found that the energy reward $R_{en}$ alone is insufficient to regularize Go1’s behavior, which is likely due to the lighter weight compared to its motor power. Thus, following the settings in [4], we add the fixed auxiliary reward $R_{aux}$. This auxiliary reward is derived mainly from safety concerns, such as penalizing limb-ground collision, out-of-range joint position, and high frequency joint action. The details of $R_{aux}$ can be found on the project website. Compared to [4], we did not include any gait-related rewards.
IV-A2 Curriculum and Domain Randomization
We deployed curriculum technique on command velocities. The sampling range of linear and angular velocities start at $[-1,1]$ m/s and $[-1,1]$ rad/s. The sampling range increases when the total reward achieves a certain threshold, and the maximal sampling range is set at $[-5,5]$ m/s and $[-5,5]$ rad/s. The ground coefficient of friction is randomized between $[0.05,1.5]$. We also added a disturbance to the robot mass with a uniform random value in $[-0.1,3.0]$ kg. A uniformly distributed noise was added to the observation. All these randomized domain parameters are renewed every time the training episode resets, except observation noise is resampled after every time step.
IV-A3 Amendments for Training on Terrains
All previous techniques are deployed on flat ground trainings. We also tested adding energy regularization while training on terrains. The training terrain shape is rough slope adapted from [6].
IV-B Translation and Rotation
Figure 4: Gait under different command velocities when $\alpha_{en}=0.0$. The policy shows bouncing gait across all command velocities, which is not an energy-efficient choice of gait. We also demonstrate snapshots of bouncing gait, which is the four legs touches and leaves the ground almost simultaneously.
Figure 5: Gait comparison of ANYmal-C between energy regularization and original legged gym policy [6]. Similar gait transition from walking to trotting also appears when energy regularization is applied. In the original legged gym policy, the lifting height of rear right leg is very low, so it has several unexpected mild contacts with the ground. Videos of ANYmal-C simulation can be found on our project website.
@@ IV-B Translation and Rotation
To demonstrate natural emergence of efficient locomotion gaits, we evaluate the trained walking policy under different reference velocities. Fig. 2 demonstrates a trial run where the legged robot was commanded to move from $\hat{v}_{x}=0.0$, to $2.5$ m/s. The robot accelerates at $0.5m/s^{2}$. We plot the gait recorded from the Go1 real-world deployment. We can see that our trained policy exhibits four-beat walking at low speed (around $0.0$ to $0.4$ m/s), where four legs touches the ground in front-left, front-right, rear-left, rear-right sequence. Then, the robot shows two-beat walking (around $0.4$ to $1.1$ m/s) where the four legs touch the ground in diagonal pairs and present observable moments where four legs touches the ground at the same time. At medium speed (around $1.1$ to $1.7$ m/s), the policy exhibits an trotting gait. At this gait, the four legs still touch the ground in diagonal pairs, but there is neither noticeable moment where the four legs touch the ground at the same time, nor moment where all four legs are in the air. At high speed (around $1.7$ m/s and beyond), the trained policy exhibits a fly-trotting gait, which is similar to trotting gait, except there are observable moments where all four legs are in the air. This gait transition is endorsed by previous works [2, 5] that walking and trotting are respectively the most energy-efficient gaits under low and high speeds.
The result in Fig. 2 is generated with $\alpha_{en}=1.0$. In Fig. 3(a) and Fig. 3(b), we compare CoT and linear velocity tracking results across different $\alpha_{en}$. When weight is small (like $0.0$ or $0.5$), the effect of energy regularization is minor, so the generated policy has a much higher CoT. Fig. 4 shows that when $\alpha_{en}=0.0$, the robot exhibits a bouncing gait across the whole velocity domain, which is not efficient [2].
The proposed energy regularization can be similarly applied to other quadruped platforms. Experiments were conducted on the ANYmal-C simulation environment [6] with the same scaling constants $\sigma_{en,x}=1000$ and $\sigma_{en,z}=500$ as well as the regularization weight $\alpha_{en}=1.0$. The generated policy similarly showed preferred gait transition from walking to trotting, which is show in Fig. 5. As shown in Fig. 3(d), it also successfully reduced CoT in comparison to the original legged gym settings [6].
IV-C Circle Tracking
(a) Semicircle Tracking Sketch Plot
(b) Semicircle Tracking Error Accumulation
Figure 6: Open-loop semicircle tracking comparison between $\alpha_{en}=0.0$ and $\alpha_{en}=1.0$. Since energy regularization reduces random motions irrelevant to target motions, the tracking error for $\alpha_{en}=1.0$ accumulates considerably slower than $\alpha_{en}=0.0$.
We also tested the circle-tracking capacity of the policy. We ask the robot to walk with linear velocity $1.0$ m/s and angular velocity $0.5$ rad/s. This is equivalent to asking the robot to follow a circle with radius $2$ m. We moved the robot in a semicircle with predefined center in simulation and recorded its position at each time step. Assume the distance between the robot and the circle center is $d(t)$ at time step $t$, the error $e(t)$ is defined as the absolute value of the difference between $d$ and the $2$ m pre-defined radius, $|d(t)-2|$. We compare the accumulated error $\sum_{i=1}^{T}e(t)$ between $\alpha_{en}=0.0$ and $1.0$ policies.
The results are demonstrated in Fig. 6. The trajectory radius of $\alpha_{en}=1.0$ policy is closer to the $2$ m target and the error also accumulates much slower. This is because that energy regularization can effectively reduce motions that are not directly related to the target $1.0$ m/s, $0.5$ rad/s motion.
IV-D Terrain Clearance
The effect of energy regularization on quadruped gait can also be deployed on terrain clearance. Since it is hard to walk either very fast or very slow on a rough slope, the training velocity range is limited to only $[-1.5,1.5]$ m/s. When no energy regularization presents, the quadruped robot also tends to show a bouncing gait. When $\alpha_{en}=1.0$, a more natural trotting gait appears. As indicated in Fig. 3, the policy generated from $\alpha_{en}=1.0$ is also a more energy efficient gait.
We tested the trained policy on the real Go1 quadruped robot. Fig. 7 shows experiment snapshots. The robot successfully cleared a 20 cm step covered by paper boards, which is a considerable high step compared to the size of Go1.
(a) Snapshots of the quadruped robot climbing over randomly distributed paper boards.
(b) Snapshots of the quadruped robot climbing over a 20 cm step covered by paper boards.
Figure 7: Quadruped robot clearing terrains.
V Discussion and Conclusion
@@
This paper presented a novel approach to energy-efficient locomotion in quadruped robots by implementing a simplified, energy-centric reward strategy within a reinforcement learning framework. Our method demonstrated that quadruped robots, specifically Unitree Go1, could autonomously develop and transition between various gaits across different linear and angular velocites without relying on predefined gait patterns or intricate reward designs. The adaptive energy reward function, adjusted based on velocity, enabled these robots to select the most energy-efficient locomotion strategies naturally.
There are more extended locomotion tasks that we can make use of the energy-centric training, where the motion reward is no longer velocity tracking but includes more task-specific rewards. This would require the system to dynamically adjust its strategy in response to different task rewards. While this study concentrated on locomotion tasks, the underlying principle of leveraging energy efficiency to drive behavior selection holds broader potential. Future work could explore applying this energy-centric approach across different robotic tasks. For instance, manipulation and interaction tasks could also benefit from strategies prioritizing energy efficiency, potentially finding natural, efficient behaviors analogous to those observed in biological systems [34, 35]. Such a framework would align robotic systems more closely with sustainability principles and environmental consciousness.